# Probabilistic Forecasting with Coherent Aggregation

### v3

@@ Appendix H Efficient Estimation of Scoring Rules
Figure 10: Training and validation Continuous Ranked Probability Score curves on the SF Bay Area Traffic dataset. We show curves for DeepCoFactor varying as a function of the gradient steps. Each curve uses a different number of Monte Carlo samples to estimate CRPS.
@@
$\hat{\mathrm{CRPS}}\left(y,Y\right)=\frac{1}{N}\sum^{N}_{i=1}\,|y_{i}-y|-\frac{1}{2N(N-1)}\sum^{N}_{i,j=1}|y_{i}-y_{j}|$ (26)
@@
Where $y_{i},y_{j}\quad i,j=1,...,N$ are independent samples of the random variable $Y$, and $y$ a possible observation.
@@
To assess the effect of sample size on the CRPS Monte Carlo estimator, we ran diagnostics throughout the model training process. Figure 10 shows the training and validation loss trajectories for varying Monte Carlo sample sizes, using model configurations from Table 4. For reference, we include the analytical CRPS of truncated normals [46], despite minor distributional differences compared to Eqn. 10. For a comprehensive review of alternative CRPS estimators we refer to Zamo & Naveau [55].
@@
We found that reducing the number of Monte Carlo samples had minimal impact on both training and validation scores. Given GPU memory constraints, decreasing the sample size effectively reduces computational overhead without significantly affecting model performance. We observed similar phenomenons with the energy score learning objective, and its relationship to Monte Carlo samples.

### v4

@@ Appendix H Efficient Estimation of Scoring Rules
Figure 10: Training and validation Continuous Ranked Probability Score curves on the SF Bay Area Traffic dataset. We show curves for CLOVER varying as a function of the gradient steps. Each curve uses a different number of Monte Carlo samples to estimate CRPS.
@@
$\hat{\mathrm{CRPS}}\left(y,Y({\bm{\theta}})\right)=\frac{1}{N}\sum^{N}_{i=1}\,|y_{i}-y|-\frac{1}{2N(N-1)}\sum^{N}_{i,j=1}|y_{i}-y_{j}|$ (26)
@@
Where $y_{i},y_{j}\quad i,j=1,...,N$ are independent samples of the random variable $Y({\bm{\theta}})$, and $y$ a possible observation.
@@
To assess the effect of sample size on the CRPS Monte Carlo estimator, we ran diagnostics throughout the model training process. Figure 10 shows the training and validation loss trajectories for varying Monte Carlo sample sizes, using model configurations from Table 4. For reference, we include the analytical CRPS of truncated normals [48], despite minor distributional differences compared to Eqn. 10. For a comprehensive review of alternative CRPS estimators, we refer to Zamo & Naveau [57].
@@
We found that reducing the number of Monte Carlo samples had minimal impact on both training and validation scores. Given GPU memory constraints, decreasing the sample size effectively reduces computational overhead without significantly affecting model performance. We observed similar phenomena with the energy score learning objective and its relationship to Monte Carlo samples.
Appendix I Note on the Reparameterization Trick
In this section, we discuss the importance of the reparameterization trick in the context of hierarchical series and our multivariate factor model. We explain how it enables us to maintain the network’s differentiability while optimizing arbitrary learning objectives, regardless of the complexity of our probabilistic model.
Let a disaggregated series be
$\hat{Y}({\bm{\theta}})=\mu({\bm{\theta}})+\mathrm{Diag}(\sigma({\bm{\theta}}))\mathbf{z}+\mathbf{F}({\bm{\theta}}){\bm{\epsilon}}\in\mathbb{R}^{N_{b}}.$ (27)
Let the hierarchical series under aggregation constraints be
$\tilde{Y}({\bm{\theta}})=\mathbf{S}_{[i][b]}\hat{Y}({\bm{\theta}})_{+},$ (28)
with ${\bm{\epsilon}}\sim N(\mathbf{0},\mathrm{I}_{N_{k}})$ and $\mathbf{z}\sim N(\mathbf{0},\mathrm{I}_{N_{b}})$ the reparameterization random variables. For simplicity of exposition, we consider a single entry in the reconciled multivariate series $\tilde{Y}({\bm{\theta}})$ that we denote $Y({\bm{\theta}})=\mathbf{S}_{i,[b]}\hat{Y}({\bm{\theta}})_{+}$.
For a differentiable learning objective $\mathcal{L}:\mathbb{R}^{|{\bm{\theta}}|}\mapsto\mathbb{R}$, the loss gradient with respect to the network parameters ${\bm{\theta}}$ can be obtained as
$\nabla_{{\bm{\theta}}}\mathbb{E}_{Y}\left[\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\right]$ $=\nabla_{{\bm{\theta}}}\left[\int_{Y}\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\mathbb{P}(Y|{\bm{\theta}})\delta Y\right]$ (29)
$=\int_{Y}\nabla_{{\bm{\theta}}}\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\mathbb{P}(Y|{\bm{\theta}})\delta Y+\int_{Y}\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\nabla_{{\bm{\theta}}}\mathbb{P}(Y|{\bm{\theta}})\delta Y$
$=\mathbb{E}_{Y}\left[\nabla_{{\bm{\theta}}}\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\right]+\int_{Y}\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\nabla_{{\bm{\theta}}}\mathbb{P}(Y|{\bm{\theta}})\delta Y.$
Note that the expectation’s gradient is not the expectation of the gradient due to the second term that captures the dependence of the probability of $Y({\bm{\theta}})$ on the parameters of the network. The second term may face challenges when the probability of $Y({\bm{\theta}})$ does not have an analytical form or its gradient $\nabla_{{\bm{\theta}}}\mathbb{P}(Y|{\bm{\theta}})$ poses challenges for automatic differentiation tools. The solution proposed by Kingma & Welling [29] reparameterizes the target variable to recover the differentiability of the gradient loss in Eqn. 29.
Using the chain rule and the reparameterization trick, the loss’ gradient can be recovered:
$\nabla_{{\bm{\theta}}}\mathbb{E}_{Y}\left[\mathcal{L}\left(\,Y({\bm{\theta}})\,\right)\right]$ $=\nabla_{{\bm{\theta}}}\mathbb{E}_{{\bm{\epsilon}},\mathbf{z}}\left[\mathcal{L}\left(\mathbf{S}_{[i][b]}\hat{Y}({\bm{\theta}})\right)\right]$ (30)
$=\mathbb{E}_{\mathbf{z},{\bm{\epsilon}}}\left[\nabla_{{\bm{\theta}}}\mathcal{L}\left(\mathbf{S}_{i,[b]}\hat{Y}({\bm{\theta}})\right)\right]$
$=\mathbb{E}_{\mathbf{z},{\bm{\epsilon}}}\left[\frac{\delta\mathcal{L}}{\delta Y}\cdot\mathbf{S}_{i,[b]}\cdot\left(\nabla_{{\bm{\theta}}}\hat{Y}\left({\bm{\theta}}\right)\right)\right]$
$=\mathbb{E}_{\mathbf{z},{\bm{\epsilon}}}\left[\frac{\delta\mathcal{L}}{\delta Y}\cdot\mathbf{S}_{i,[b]}\cdot\left[\nabla_{{\bm{\theta}}}\mu+\nabla_{{\bm{\theta}}}\;\sigma\;\mathbf{z}+\nabla_{{\bm{\theta}}}\mathbf{F}{\bm{\epsilon}}\right]_{+}\right].$
As highlighted in Section 5, the key insight from Eqn. 30 is that the reparameterization trick provides a highly flexible framework. It is compatible with any differentiable learning objective and can accommodate differentiable constraints that extend beyond traditional aggregation constraints, such as those in Eqn. 28. Furthermore, the techniques explored in this paper can be generalized to continuous distributions [13, 45, 24] beyond Gaussian random variables, opening up exciting avenues for future research.